OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1)

OpenStudy (anonymous):

you can use any method you like, including checking the answers to see which one works

OpenStudy (theslytherinhelper):

Alright, so, do you know the first step?

OpenStudy (anonymous):

this must be a different problem right?

OpenStudy (theslytherinhelper):

Uh...Those aren't even the equations you're using. x - 4y = 1 2x - 9y = 3 First, choose a variable. Either y or x

OpenStudy (anonymous):

oops sorry.

OpenStudy (anonymous):

what problem do you need to solve? is it this one \[x - 4y = 1\]\[ 2x - 9y = 3\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can solve the first equation for \(x\) via \(x=4y+1\) and then substitute it in to the second equation to get \[2(4y+1)-9y=3\]

OpenStudy (anonymous):

\[x-4y=1 <=> y=x/4 -1/4\] put y into the second equation \[2x - 9y = 2x - 9(x/4 - 1/4) = 3 <=> x=-3\] thus \[y=x/4 -1/4 = -1\]

OpenStudy (anonymous):

then solve that one in a couple of steps multiply out to get \[8y+2-9y=3\] combine like terms to get \[-y+2=3\] subtract 2 get \[-y=1\] so \[y=-1\]

OpenStudy (anonymous):

so the first part is -1 ?

OpenStudy (anonymous):

it is \(y=-1\) then you can solve for \(x\)

OpenStudy (anonymous):

or, since you have only one choice with the second number as \(-1\) you can assume that it is the choice \((-3,-1)\)

OpenStudy (anonymous):

no :/ the teacher cant really explain this too good

OpenStudy (anonymous):

u jus said the first part would be -1 so how would it be second ?

OpenStudy (theslytherinhelper):

That's the y. (x, y)

OpenStudy (anonymous):

@mikaa_toxica13 i solved for \(y\) not for \(x\) \(y\) is the second coordinate, not the first coordinate

OpenStudy (anonymous):

oh okay i see

OpenStudy (anonymous):

sorry i got confused

OpenStudy (anonymous):

now that you know \\(y=-1\) and you know \(x - 4y = 1\) that means \[x-4(-1)=1\] so \[x+4=1\iff x=-3\]

OpenStudy (anonymous):

first coordinate is \(-3\) and the second is \(-1\)

OpenStudy (anonymous):

okay thank u soo much !!